OpenStudy (anonymous):

Trying to solve this one, 20x2+36x-8=0 ??? I found 2 solutions: x = -2 x = 1/5 I pulled out the like factors: 20x2 + 36x - 8 = 4 • (5x2 + 9x - 2) Split the middle term: 5x2 + 9x - 2 then...-10 + 1 = -9 -5 + 2 = -3 -2 + 5 = 3 -1 + 10 = 9 Did I do this wrong? I'M SOOOO CONFUSED!!!

OpenStudy (cj49):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

I rewrote the polynomial splitting the middle term using the two factors found in question, -1 and 105x2 - 1x + 10x - 2...CORRECT?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Okay...

OpenStudy (cj49):

x = \[9\pm \sqrt{121}/10\] x = [ 9 ± 11 ) ] / 10 x = 2 , x = - 1/5 OR (5x + 1)(x - 2) = 0 x = - 1 / 5 , x = 2

OpenStudy (paki):

for factoring quadratic formula is the best....

OpenStudy (unklerhaukus):

you can check your solution by plugging them into the original equation, if you get 0 = 0 your solutions solve the equation